OpenStudy (anonymous):

The sum of two consecutive integers is at least 36. What is the least possible pair of integers? A.19 and 20 B.17 and 18 C.18 and 19 D.16 and 17

OpenStudy (anonymous):

i tired doing a quiz on this but i failed

OpenStudy (markchernioglo):

oh

OpenStudy (markchernioglo):

well what did you put?

OpenStudy (maddy1251):

What did you put?

OpenStudy (maddy1251):

Which pair(s) of these numbers add up to 36?

OpenStudy (markchernioglo):

yup

OpenStudy (anonymous):

I put C cause idk

OpenStudy (maddy1251):

A.19 and 20 B.17 and 18 C.18 and 19 D.16 and 17 We will add these.

OpenStudy (markchernioglo):

So try all your pairs: 16+17 = 33 17+18 = 35 18+19 = 37 19+20 = 39 So The first two ones don't add up to 36, so they are out. Next you are left with 18+19=37 and 19+20=39 So 37<39 That means 18+19 < 19+20 So the smallest pair of integers that add up to at least 36 is 18+19.

OpenStudy (maddy1251):

@markchernioglo you beat me to it!

OpenStudy (markchernioglo):

:)

OpenStudy (markchernioglo):

lol

OpenStudy (maddy1251):

Hey I only have one hand to type with, soo unfair :p

OpenStudy (markchernioglo):

its alright maddy ill give you a medal

OpenStudy (maddy1251):

I will medal you too!

OpenStudy (markchernioglo):

you tried

OpenStudy (maddy1251):

You had a great explanation. And I did try :p

OpenStudy (markchernioglo):

lol

OpenStudy (anonymous):

ok ok so all u did was see wat was greater

OpenStudy (maddy1251):

I was typing out my answer, and poof you put yours in.

OpenStudy (maddy1251):

Not greater, but clsoer to 36 as we could get.

OpenStudy (markchernioglo):

no what added to 36

OpenStudy (markchernioglo):

yup maddy is correct

OpenStudy (maddy1251):

None added to 36, but we did have 37, which is close enough, because the rest where greater than 36.

OpenStudy (markchernioglo):

lol good job everyone

OpenStudy (maddy1251):

See, I am good! :D Haha.

OpenStudy (markchernioglo):

yup close to 36

OpenStudy (markchernioglo):

lol good job

OpenStudy (anonymous):

thank you both

OpenStudy (markchernioglo):

np

OpenStudy (maddy1251):

In math, we don't always get handed easy situations, like this, so we gotta find what works BEST. Not a problem. :)

OpenStudy (markchernioglo):

anytime

OpenStudy (anonymous):

can yall be my math teachers